OpenStudy (anonymous):

What is the distance between the points (80, 55) and (20, 44)?

OpenStudy (anonymous):

\[DISTANCE = \sqrt{(X _{1} - X _{2})²+(Y _{1} - Y _{2})²}\] X1 = 80 Y1 = 55 X2 = 20 Y2 = 44

OpenStudy (anonymous):

I did that but my answer was way to high.

OpenStudy (anonymous):

sqrt((80-50)² + (55-44)²) sqrt(30² + 11²) sqrt(900 +121) = 31.953

OpenStudy (anonymous):

is the correct answer 11?

Directrix (directrix):

I used the online distance between 2 points calculator and got 61 as the distance. http://www.easycalculation.com/analytical/distance.php From above: sqrt((80-50)² + (55-44)²) --> should be sqrt(80-20)^2 + (55-44)² sqrt(60² + 11²) sqrt(3600 +121) sqrt(3721) = 61